OpenStudy (anonymous):

Which ordered pair is a solution of the system: 3x + 2y = 4 and -x + 3y = -5 ?

OpenStudy (fibonaccichick666):

So first, can you try and solve both eq for the same variable please?

OpenStudy (anonymous):

Or just solve one equation for one variable. Perhaps a more efficient version.

OpenStudy (anonymous):

i dont know how thats the problem lol

OpenStudy (anonymous):

Can you solve this equation for x? 3y - x = 5 It doesn't matter if the y is still there.

OpenStudy (anonymous):

oh yea okay, let me solve it real quick

OpenStudy (fibonaccichick666):

ok, then do the same for the other

OpenStudy (anonymous):

okay so for 3y-x=5 so i got y=2

OpenStudy (anonymous):

and the other one is -1 , am i right?

OpenStudy (fibonaccichick666):

uhm, you shouldn't get a number. You should get an eq set up like this x=blah blah y blah or y=blah x blah blah

OpenStudy (mathmale):

I'd like to suggest a different method of solution: elimination by addition/subtraction. 3x + 2y = 4 and -x + 3y = -5 can be re-written as 3x + 2y = 4 -x + 3y = -5 Multiply the second equation by 3. Then add this result to 3x + 2y = 4. What happens?

OpenStudy (fibonaccichick666):

ahh, the matrix way. I totally forgot they do that at this level Nice addition @mathmale